The density of a 1.96 M solution of LiBr in acetonitrile (CH3CN) in 0.826 g/mL. Calculate the concentration of this solution in (a) molality, (b) mole fraction of LiBr, and (c) mass percent of CH3CN.
Answer #1

Assume the volume of the solution is 1L or 1000 mL

Mass of solution = density x Volume = 0.826 g/mL x 1000 = 826g

Moles of LiBr in the solution = Molarity x Volume (in litres)

= 1.96 x 1 = 1.9mol

Mass of LiBr in the solution = moles x molar mass

= 1.96 mol x 86.9 g/mol = 170.324 g

Mass of acetonitrile. = Total mass - mass of LiBr

= 826 - 170.324 = 655.676 g

Moles of acetonitrile = mass / molar mass = 655.676 / 41 = 15.99 mol

(a) molality = moles solute / kg solvent = 1.96/0.655676 = 3m (approx)

(b) moles fraction of LiBr = moles of LiBr / total moles

= 1.96 / ( 1.96 + 15.99 ) = 0.11

(c) mass percent of CH3CN

= (Mass of acetonitrile/total mass of solution )*100%

= (655.676 / 826)*100 = 79.4%